OpenStudy (anonymous):

A circle is centered at (3, 8) and is tangent to the x-axis. Write its equation in standard form.

OpenStudy (anonymous):

i just need help not the answer

OpenStudy (jdoe0001):

 we know the center of it, so we just need the radius notice we are given 2 points already... so what's the distance between (3, 8) and (3, 0) ? well \(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(3\quad ,&8)\quad &(3\quad ,&0) \end{array}\qquad d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\)

OpenStudy (jdoe0001):

"and is tangent to the x-axis" <-- meaning is touching it, thus

OpenStudy (anonymous):

i got (x - 3)^2 + (y - 8)^2 = 64

OpenStudy (anonymous):

is that correct?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

tysm !

OpenStudy (jdoe0001):

yw